Cash Disbursement
Timber Company is in the process of preparing its budget for next year. Cost of goods sold has been estimated at 70 percent of sales. Lumber purchases and payments are to be made during the month preceding the month of sale. Wages are estimated at 15 percent of sales and are paid during the month of sale. Other operating costs amounting to 10 percent of sales are to be paid in the month following the month of sale. Additionally, a monthly lease payment of $14,000 is paid for computer services. Sales revenue is forecast as follows

Month Sales Revenue
February $170,000
March 210,000
April 220,000
May 260,000
June 240,000
July 280,000

Required
Prepare a schedule of cash disbursements for April, May, and June.
Do not use a negative sign with your answers.

Timber Company
Schedule of Cash Disbursements
April, May, and June
April May June
Lumbers purchases
Wages
Operating expenses
Lease payment
Total disbursements

Classify each of the following as structural, organizational, or activity cost drivers.

(a) Apple Inc. reorganizes production facilities from a layout in which all similar types of machines are grouped together to a layout in which a set of machines is designated for the production of a particular product and that set of machines is grouped together.

Structural

Organizational

Activity


(b) A cable television company decides to start offering telephone service.

Structural

Organizational

Activity


(c) IBM Corporation decides to stop making personal computers.

Structural

Organizational

Activity


(d) Canon decides to start making high-volume photocopy equipment to compete head-to-head with Xerox.

Structural

Organizational

Activity


(e) The number of meals a cafeteria serves.

Structural

Organizational

Activity


(f) The number of miles a taxi is driven.

Structural

Organizational

Activity


(g) A company eliminates the position of supervisor and has each work group elect a team leader.

Structural

Organizational

Activity


(h) Tesla empowers employees to halt production if a quality problem is identified.

Structural

Organizational

Activity


(i) The number of tons of grain a ship loads.

Structural

Organizational

Activity


(j) Northbrook Mall decides to build space for 80 additional stores.

Structural

Organizational

Activity
Answer #1

Ans- Timber Company

Schedule of Cash Disbursements

April, May and June

April May June
Lumbers Purchases $182,000 $168,000 $196,000
Wages $33,000 $39,000 $36,000
Operating Expenses $21,000 $22,000 $26,000
Lease Payment $14,000 $14,000 $14,000
Total Disbursements $250,000 $243,000 $272,000

Working Note:-

Lumber Purchases 260,000*70%= $182,000 240,000*70%=$168,000 280,000*70%=$196,000
Wages 220,000*15%=$33,000 260,000*15%= $39,000 240,000*15%= $36,000
Operating Expenses 210,000*10%= $21,000 220,000*10%= $22,000 260,000*10%= $26,000
(a) Organizational
(b) Structural
(c) Structural
(d) Structural
(e) Activity
(f) Activity
(g) Organizational
(H) Organizational
(i) Activity
(j) Structural

Kindly give thumbsup if u like my answer...Thanks!!!